# A fully Bayesian approach for the imputation and analysis of derived outcome variables with missingness

### v1

@@ 5 Discussion
The advantages of modelling the source variables directly include the ability to incorporate informed priors specific to the individual source variables and model missingness within a fully Bayesian model following our proposed method. One could also incorporate adjustments for other sources of bias beyond missingness such as preferential sampling (Campbell et al. 2022) and measurement error. This is an important future goal for the microcephaly model given the issues considered by Harville et al. (2020). Including additional variables in the model (e.g., variables that are potentially prognostic factors for microcephaly including exposure to teratogens, and maternal age (Liu et al. 2019)) is another important objective.
There will also be disadvantages to modelling the source variables directly. The additional complexity involved compared to fitting a simpler model involving only the derived variable will likely increase the risk of model misspecification. This includes the risk of misspecifying the priors. Indeed one must be careful to ensure that the “implied” prior on the distribution of the derived variable is reasonable and this is not always obvious (Lockhart 2019). In the microcephaly example, much effort was involved in selecting priors for the BsNmN model such that the implied prior distribution on the risk of microcephaly was reasonable. (The $\textrm{Beta}(1.26,2.39)$ prior for the Bernoulli model was selected in order for the two models to have (implied) priors matching as closely as possible; see Figure 4).
Finally, we note that there is also a risk of model misspecification when using MICE. In the microcephaly example, the Bernoulli model with the entire dataset analysis appears to be biased ($\hat{\theta}=12.73\%$ vs the truth $\theta=11.69\%$). We suspect that this is due to the distributional assumptions invoked in the MICE algorithm being too simplistic for these data. Specifically, a normal distribution was assumed for missing gestational age values when, in reality, the data are highly skewed. Figure 7 shows how this results in a slight overestimation of the missing $Z_{2}$ values, which, in turn, leads to a relatively large overestimation in the number of microcephaly cases.

### v2

@@ 5 Discussion
If the amount of missingness is very small or negligible, applying a missing value procedure should not drastically change one’s estimates, even if this procedure involves a misspecified model for the missing values. Within Gelman’s three step method, if the amount of missingness is negligible, an incorrect imputation model should have only a minimal negative effect. In our proposed fully-integrated Bayesian approach, the model for the missing values and the outcome model are one and the same. Therefore if this model is misspecified, the consequences might be more substantial even if there is negligible missingness. This model might also be more complex than the one originally considered for analysis. For instance, a researcher might be able to make relatively simple assumptions about the distribution of the derived outcome variable (e.g., be able to assume that microcephaly status is binomial), but have difficulty in making reasonable distributional assumptions for the source variables (e.g., what are reasonable distributions to assume for gestational age and head circumference?).
The advantages of modeling the source variables directly include the ability to incorporate informed priors specific to the individual source variables and model missingness within a fully Bayesian model following our proposed method. One could also incorporate adjustments for other sources of bias beyond missingness such as preferential sampling (Campbell et al. 2022) and measurement error. This is an important future goal for the microcephaly model given the issues considered by Harville et al. (2020).
Finally, we note that our proposed approach allows one to conduct the analysis of data with missingness within an entirely Bayesian framework. In some instances, one might wish to fit a Bayesian imputation model for the missing values and frequentist model for the main analysis (i.e., for the outcome model). In such a case, Rubin’s rules could be appropriately and easily applied. Alternatively, if both the imputation model and the outcome model of interest must be frequentist, von Hippel & Bartlett’s rules (Von Hippel & Bartlett 2021) may be more appropriate. To the best of our knowledge, if the imputation model is frequentist and the outcome model is Bayesian (an admittedly obscure scenario), there does not appear to be a good approach for obtaining estimates.